OpenStudy (anonymous):

sin 2 theta if sin theta = 3/5 and theta is the first quadrant

OpenStudy (debbieg):

\[\sin(2\theta)=2\sin \theta \cos\theta\] and by the fundamental identity \[\sin^2\theta + \cos ^2 \theta=1\] You have sinΘ. And using the Fundamental Identity, along with the information about the quadrant, you can obtain cosΘ. Plug these values back into the double angle identity, and you will get \[\sin(2\theta)\]

OpenStudy (anonymous):

thank you